# Residually Constructible Extensions

### v4

@@ Abstract.
Let $T$ be an o-minimal theory expanding $\mathrm{RCF}$ and $T_{\mathrm{convex}}$ be the common theory of its models expanded by predicate for a non-trivial $T$-convex valuation ring. We call an elementary extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ res-constructible if there is a tuple $\overline{s}$ in $\mathcal{O}_{*}$ such that the projection $\operatorname{\mathbf{res}}(\overline{s})$ of $\overline{s}$ in the residue field sort is $\mathrm{dcl}$-independent over the residue field $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$ of $(\mathbb{E},\mathcal{O})$ and such that $\operatorname{\mathrm{dcl}}(\mathbb{E},\overline{s})=\mathbb{E}_{*}$. We study factorization properties of res-constructible extensions. Our main result is that a res-constructible extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ has the property that all $(\mathbb{E}_{1},\mathcal{O}_{1})$ with $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ are res-constructible over $(\mathbb{E},\mathcal{O})$, if and only if $\mathbb{E}_{*}$ has countable dimension over $\mathbb{E}$ or the value group $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{*})$ contains no uncountable well-ordered subset. This analysis entails complete answers to [9, Problem 5.12].
@@
The second author’s work was supported by EPSRC DTP 2224 [EP/W524372/1]
@@ 1. Introduction
By results of van den Dries and Lewenberg [3, (3.10), (3.13), (3.14)], the common theory $T_{\mathrm{convex}}$ in the language $L_{\mathrm{convex}}:=L\cup\{\mathcal{O}\}$ of all models of $T$ expanded by a unary predicate $\mathcal{O}$ interpreted as a non-trivial $T$-convex valuation ring, is complete and weakly o-minimal.
@@
Subsequent to proving Theorem A, in Section  4, we will turn to [9, Problem 5.12], and give a complete answer, dependent on the properties of $\mathbb{E}_{*}$. We will also show that in the power-bounded setting, this dependence can be reduced to considering properties of $\mathbb{E}$.
@@ 3.1. Res-constructible extensions
Let $\mathbb{E}\preceq\mathbb{E}_{*}\models T$ and let $\mathcal{O}_{*}$ be a non-trivial $T$-convex subring of $\mathbb{E}_{*}$. A possibly infinite tuple $\overline{y}:=(y_{i}:i\in I)$ in $\mathcal{O}_{*}$ will be called a $\mathcal{O}_{*}$-res-construction over $\mathbb{E}$ if $\big(\operatorname{\mathbf{res}}_{\mathcal{O}_{*}}(y_{i}):i\in I\big)$ is $\operatorname{\mathrm{dcl}}_{T}$-independent over $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$. We then call the extension $\mathbb{E}\langle y_{i}:i\in I\rangle$ of $\mathbb{E}$ $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$ (by the res-construction $(y_{i}:i\in I)$). When $\mathcal{O}_{*}$ is clear from the context we will just say res-construction and res-constructible. We call an extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ residually constructible (res-constructible for short) when $\mathbb{E}_{*}$ is $\mathcal{O}_{*}$-res constructible over $\mathbb{E}$.
@@ 3.2. Right factors of res-constructible extensions
Suppose that $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{1})$ contains no uncountable ordinal and $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$. Then for every $\operatorname{\mathrm{dcl}}_{T}$-basis $B$ of $\mathbb{E}_{1}$ over $\mathbb{E}\langle\overline{x}\rangle$, there is an ordinal-indexing $(b_{i}:i<\lambda)$ of $B$, such that for all $i<\lambda$, $\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},b_{j}:j<i\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},b_{j}:j\leq i\rangle)$.
@@
By the hypothesis we can find $(s_{i}:i<\aleph_{1})$, $r$ and $(r_{i}:i<\aleph_{1})$ in $\mathbb{E}_{*}$ such that $\operatorname{\mathbf{res}}(r)$, $\operatorname{\mathbf{res}}(r_{i})_{i<\aleph_{1}}$, $\operatorname{\mathbf{res}}(s_{i})_{i<\kappa}$ are a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\operatorname{\mathbf{res}}(\mathbb{E}_{*})$ over $\operatorname{\mathbf{res}}(\mathbb{E})$ and moreover $\mathbb{E}_{0}:=\mathbb{E}\langle s_{i}:i<\kappa\rangle$ contains a sequence $(x_{i}:i<\aleph_{1})$ such that $\operatorname{\mathbf{v}}(x_{i})>\operatorname{\mathbf{v}}(x_{j})$ for all $i<j$.
Set $\mathbb{F}_{i}^{\prime}:=\mathbb{E}_{0}\langle r+r_{j}x_{j}:j<i\rangle$. We claim that $\mathbb{F}:=\mathbb{F}^{\prime}_{\aleph_{1}}$ is not res-constructible. Suppose toward contradiction that it is: then there is a sequence $(\mathbb{F}_{i})_{i<\aleph_{1}}$, such that
@@ 4. Application to Tressl’s pseudo-completions
We first observe that if $\mathbb{K}\prec\mathbb{K}_{1}\models T$, $V$ is a non-trivial $T$-convex valuation ring of $\mathbb{K}_{1}$, and $\mathbb{K}_{1}$ is $V$-res-constructible over $\mathbb{K}$, then $\mathbb{K}$ is cofinal in $\mathbb{K}_{1}$. In fact if there was $t\in\mathbb{K}_{1}$ such that $t>\mathbb{K}$, then $t$ is $V$-purely valuational over $\mathbb{K}$ so by Fact 12 $t$ is not $V$-residually generated, contradicting Lemma 20.
Now let $\overline{r}$ be an $\mathcal{O}$-res-construction of $\mathbb{E}_{1}$ over $\mathbb{E}$, then by Lemma 23, it is also a $\mathcal{O}^{\prime}$-res construction and $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\overline{r})$ is a $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res-construction for $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ over $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. In particular $\mathbb{E}_{1}$ is $\mathcal{O}^{\prime}$-res-constructible over $\mathbb{E}$ and $\mathbb{K}_{1}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ is $V:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res-constructible over $\mathbb{K}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. Thus to conclude we only need to show that $\mathbb{K}$ is cofinal in $\mathbb{K}_{1}$.
@@
If $\mathbb{E}_{*}$ is a $\mathcal{F}$-residually dense extension of $\mathbb{E}$ and $\mathcal{O}$ is any $T$-convex valuation ring of $\mathbb{E}$, then $\operatorname{\mathrm{CH}}_{\mathbb{E}_{*}}(\mathcal{O})$ is the unique extension of $\mathcal{O}$ to a $T$-convex valuation ring of $\mathbb{E}_{*}$.
To see this, note first that the extension is unique if and only if $\mathbb{E}_{*}$ does not realise the type $\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$, so we can reduce to the case where $\mathbb{E}\preceq\mathbb{E}_{*}$ is of finite $\operatorname{\mathrm{dcl}}_{T}$-dimension. Note next that if $\mathbb{E}\preceq\mathbb{E}_{\dagger}\preceq\mathbb{E}_{*}$, and $\mathcal{O}$ has a unique extension to $\mathbb{E}_{\dagger}$, say $\mathcal{O}_{\dagger}$, and $\mathcal{O}_{\dagger}$ has a unique extension to $\mathbb{E}_{*}$, then $\mathcal{O}$ must have a unique extension to $\mathbb{E}_{*}$. Thus, we can reduce further to the case where $\mathbb{E}\preceq\mathbb{E}_{*}$ is $1$-dimensional. In particular, we may assume $\mathbb{E}\preceq\mathbb{E}_{*}$ is $\mathcal{O}^{\prime}$-residually dense for some $\mathcal{O}^{\prime}\in\mathcal{F}$. If $\mathcal{O}^{\prime}\cap\mathbb{E}\subsetneq\mathcal{O}$, then Lemma 46 implies that the extension is $\mathcal{O}$-residually cofinal, whence the result follows. If they are equal, the result follows from Fact 12. Finally, if $\mathcal{O}^{\prime}\cap\mathbb{E}\supsetneq\mathcal{O}$, then any $\mathcal{O}^{\prime}$-residually dense extension cannot realize the cut $\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$ because its realization would be $\mathcal{O}^{\prime\prime}$-residual where $\mathcal{O}^{\prime\prime}:=\{t\in\mathbb{U}:|t|<\mathbb{E}^{>\mathcal{O}}\}\subsetneq\mathcal{O}^{\prime}$.
@@
If $\mathcal{F}$ is a family of $T$-convex valuation rings of $\mathbb{E}$, then the pseudo-completion of $\mathbb{E}$ with respect to $\mathcal{F}$ is unique up to isomorphism over $\mathbb{E}$.

### v5

@@ Abstract.
Let $T$ be an o-minimal theory expanding $\mathrm{RCF}$ and $T_{\mathrm{convex}}$ be the common theory of its models expanded by predicate for a non-trivial $T$-convex valuation ring. We call an elementary extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ res-constructible if there is a tuple $\overline{s}$ in $\mathcal{O}_{*}$ such that $\mathbb{E}_{*}=\operatorname{\mathrm{dcl}}(\mathbb{E},\overline{s})$, and the projection $\operatorname{\mathbf{res}}(\overline{s})$ of $\overline{s}$ in the residue field sort is $\mathrm{dcl}$-independent over the residue field $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$ of $(\mathbb{E},\mathcal{O})$. We study factorization properties of res-constructible extensions. Our main result is that a res-constructible extension $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ has the property that all $(\mathbb{E}_{1},\mathcal{O}_{1})$ with $(\mathbb{E},\mathcal{O})\prec(\mathbb{E}_{1},\mathcal{O}_{1})\prec(\mathbb{E}_{*},\mathcal{O}_{*})$ are res-constructible over $(\mathbb{E},\mathcal{O})$, if and only if $\mathbb{E}_{*}$ has countable dimension over $\mathbb{E}$ or the value group $\operatorname{\mathbf{v}}(\mathbb{E}_{*},\mathcal{O}_{*})$ contains no uncountable well-ordered subset. This analysis entails complete answers to [9, Problem 5.12].
@@
The second author’s work was supported by EPSRC DTP 2224 [EP/W524372/1] (at the University of Leeds)
For the purpose of open access, the authors have applied a Creative Commons Attribution ($C$ BY) licence to any Author Accepted Manuscript version arising from this submission. There is no data associated with this publication.
@@ 1. Introduction
By results of [3], the common theory $T_{\mathrm{convex}}$ in the language $L_{\mathrm{convex}}:=L\cup\{\mathcal{O}\}$ of all models of $T$ expanded by a unary predicate $\mathcal{O}$ interpreted as a non-trivial $T$-convex valuation ring, is complete and weakly o-minimal.
@@
Subsequent to proving Theorem A, in Section  4, we will turn to [9, Problem 5.12], and give a complete answer, dependent on the properties of $\mathbb{E}_{*}$. We will also show that in the power-bounded setting, this dependence can be reduced to considering properties of $\mathbb{E}$. Then, in Section  5, we will discuss alternative characterisations of res-constructibility, and construct an example showing that res-constructibility is not a ‘local’ property.
@@ 3.1. Res-constructible extensions
Let $\mathbb{E}\preceq\mathbb{E}_{*}\models T$ and let $\mathcal{O}_{*}$ be a non-trivial $T$-convex subring of $\mathbb{E}_{*}$. A possibly infinite tuple $\overline{y}:=(y_{i}:i\in I)$ in $\mathcal{O}_{*}$ will be called a $\mathcal{O}_{*}$-res-construction over $\mathbb{E}$ if $\big(\operatorname{\mathbf{res}}_{\mathcal{O}_{*}}(y_{i}):i\in I\big)$ is $\operatorname{\mathrm{dcl}}_{T}$-independent over $\operatorname{\mathbf{res}}(\mathbb{E},\mathcal{O})$. We then call the extension $\mathbb{E}\langle y_{i}:i\in I\rangle$ of $\mathbb{E}$ $\mathcal{O}_{*}$-res-constructible over $\mathbb{E}$ (by the res-construction $(y_{i}:i\in I)$). When $\mathcal{O}_{*}$ is clear from the context we will just say res-construction and res-constructible.
We call an extension $(\mathbb{E},\mathcal{O})\preceq(\mathbb{E}_{*},\mathcal{O}_{*})\models T_{\mathrm{convex}}$ residually constructible (res-constructible for short) when $\mathbb{E}_{*}$ is $\mathcal{O}_{*}$-res constructible over $\mathbb{E}$.
@@ 3.2. Right factors of res-constructible extensions
Suppose that $\overline{x}\triangleleft_{\mathbb{K}_{*}}\mathbb{E}_{1}$, $\operatorname{\mathbf{v}}_{\mathcal{O}_{*}}(\mathbb{E}_{1})$ contains no uncountable ordinal, and $\operatorname{\mathrm{st}}_{\mathbb{K}_{*}}(\mathbb{E}_{1})=\mathbb{K}_{*}$. Then for every $\operatorname{\mathrm{dcl}}_{T}$-basis $B$ of $\mathbb{E}_{1}$ over $\mathbb{E}\langle\overline{x}\rangle$, there is an ordinal-indexing $(b_{i}:i<\lambda)$ of $B$, such that for all $i<\lambda$, $\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},b_{j}:j<i\rangle)\neq\operatorname{\mathbf{res}}(\mathbb{E}\langle\overline{x},b_{j}:j\leq i\rangle)$.
@@
By the hypothesis we can find $(s_{i}:i<\aleph_{1})$, $r$ and $(r_{i}:i<\aleph_{1})$ in $\mathbb{E}_{*}$ such that $\operatorname{\mathbf{res}}(r)$, $\operatorname{\mathbf{res}}(r_{i})_{i<\aleph_{1}}$, $\operatorname{\mathbf{res}}(s_{i})_{i<\kappa}$ are a $\operatorname{\mathrm{dcl}}_{T}$-basis of $\operatorname{\mathbf{res}}(\mathbb{E}_{*})$ over $\operatorname{\mathbf{res}}(\mathbb{E})$ and moreover $\mathbb{E}_{0}:=\mathbb{E}\langle s_{i}:i<\kappa\rangle$ contains a sequence $(x_{i}:i<\aleph_{1})$ such that $\operatorname{\mathbf{v}}(x_{i})>\operatorname{\mathbf{v}}(x_{j})$ for all $i<j$. Set $\mathbb{F}_{i}^{\prime}:=\mathbb{E}_{0}\langle r+r_{j}x_{j}:j<i\rangle$. We claim that $\mathbb{F}:=\mathbb{F}^{\prime}_{\aleph_{1}}$ is not res-constructible. Suppose toward contradiction that it is: then there is a sequence $(\mathbb{F}_{i})_{i<\aleph_{1}}$, such that
@@ 4. Application to Tressl’s pseudo-completions
We first observe that if $\mathbb{K}\prec\mathbb{K}_{1}\models T$, $V$ is a non-trivial $T$-convex valuation ring of $\mathbb{K}_{1}$, and $\mathbb{K}_{1}$ is $V$-res-constructible over $\mathbb{K}$, then $\mathbb{K}$ is cofinal in $\mathbb{K}_{1}$. In fact if there was $t\in\mathbb{K}_{1}$ such that $t>\mathbb{K}$, then $t$ is $V$-purely valuational over $\mathbb{K}$ so by Fact 12 $t$ is not $V$-residually generated, contradicting Lemma 20. Now let $\overline{r}$ be an $\mathcal{O}$-res-construction of $\mathbb{E}_{1}$ over $\mathbb{E}$, then by Lemma 23, it is also a $\mathcal{O}^{\prime}$-res construction and $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\overline{r})$ is a $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res-construction for $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ over $\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. In particular $\mathbb{E}_{1}$ is $\mathcal{O}^{\prime}$-res-constructible over $\mathbb{E}$ and $\mathbb{K}_{1}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E}_{1})$ is $V:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathcal{O})$-res-constructible over $\mathbb{K}:=\operatorname{\mathbf{res}}_{\mathcal{O}^{\prime}}(\mathbb{E})$. Thus to conclude we only need to show that $\mathbb{K}$ is cofinal in $\mathbb{K}_{1}$.
@@
If $\mathbb{E}_{*}$ is a $\mathcal{F}$-residually dense extension of $\mathbb{E}$ and $\mathcal{O}$ is any $T$-convex valuation ring of $\mathbb{E}$, then $\operatorname{\mathrm{CH}}_{\mathbb{E}_{*}}(\mathcal{O})$ is the unique extension of $\mathcal{O}$ to a $T$-convex valuation ring of $\mathbb{E}_{*}$. To see this, note first that the extension is unique if and only if $\mathbb{E}_{*}$ does not realise the type $\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$, so we can reduce to the case where $\mathbb{E}\preceq\mathbb{E}_{*}$ is of finite $\operatorname{\mathrm{dcl}}_{T}$-dimension. Note next that if $\mathbb{E}\preceq\mathbb{E}_{\dagger}\preceq\mathbb{E}_{*}$, and $\mathcal{O}$ has a unique extension to $\mathbb{E}_{\dagger}$, say $\mathcal{O}_{\dagger}$, and $\mathcal{O}_{\dagger}$ has a unique extension to $\mathbb{E}_{*}$, then $\mathcal{O}$ must have a unique extension to $\mathbb{E}_{*}$. Thus, we can reduce further to the case where $\mathbb{E}\preceq\mathbb{E}_{*}$ is $1$-dimensional. In particular, we may assume $\mathbb{E}\preceq\mathbb{E}_{*}$ is $\mathcal{O}^{\prime}$-residually dense for some $\mathcal{O}^{\prime}\in\mathcal{F}$. If $\mathcal{O}^{\prime}\cap\mathbb{E}\subsetneq\mathcal{O}$, then Lemma 46 implies that the extension is $\mathcal{O}$-residually cofinal, whence the result follows. If they are equal, the result follows from Fact 12. Finally, if $\mathcal{O}^{\prime}\cap\mathbb{E}\supsetneq\mathcal{O}$, then any $\mathcal{O}^{\prime}$-residually dense extension cannot realize the cut $\mathbb{E}<x<\mathbb{E}^{>\mathcal{O}}$ because its realization would be $\mathcal{O}^{\prime\prime}$-residual where $\mathcal{O}^{\prime\prime}:=\{t\in\mathbb{U}:|t|<\mathbb{E}^{>\mathcal{O}}\}\subsetneq\mathcal{O}^{\prime}$.
@@
If $\mathcal{F}$ is a family of $T$-convex valuation rings of $\mathbb{E}$, then the pseudocompletion of $\mathbb{E}$ with respect to $\mathcal{F}$ is unique up to isomorphism over $\mathbb{E}$.